A business firm sells a good at the price of Rs 450.The firm has decided to reduce the price of goods to Rs 350. Consequently, the quantity demand for the good rose from 25000 units to 35000 units. Calculate the price elasticity of demand


1
Expert's answer
2020-12-06T17:49:10-0500

calculate the two midpoints;

Midpoint Quantity(Q)=(Q"_{1}" +Q"_{2}" )"\/" 2

Midpoint Price(P)=(P"_{1}" +P"_{2}" )"\/" 2

Therefore, Q=(25000+35000)"\/" 2=30000

P=(450+350)"\/" 2=400

Next,we calculate PED using the ordinary formula.

PED="\%" Change in Quantity(Q"_{2}" -Q"_{1}" )/Midpoint Q divided by "\%" change in Price(P2-P1)"\/" Midpoint P

=(35000-25000)"\/" 30000 divided by (350-450)"\/" 400

=0.33"\/" -0.25

=-1.32